if sinx + sin 2 x+ sin 3 x = 1 , then prove that cos 6 x- 4cos 4 x + 8cos 2 x= 4 .This is the hardest question and I ...

if sinx+sin2x+sin3x=1then prove that cos6x-4cos4x+8cos2x= 4 .This is the hardest question and I ...

Grade:10

1 Answers

kanha kasniya
7 years ago
Sin(x) +sin^3(x) = 1-sin^2(x) => sin^2(x)[1+sin^2(x)]^2=cos^4(x)=> [1-cos^2(x)][2-cos^2(x)]^2=cos^4(x)Cos^6(x) - 4cos^4(x) + 8cos^2(x) = 4